This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for all experts' activities  NA61/SHINE logo
Message ID: 371     Entry time: Wed, 11.03.2026 21:40:55
Author: B. Lyon 
Subject: TPC HV settings for 2026 p+p run 
VTPC-1 Gap TPC VTPC-2 FTPC 1 MTPC-L MTPC-R FTPC 2 FTPC 3
Wed, 11.03.2026 20:50:00 500V/13kV 500V/8.16kV 500V/13kV 500V/17.3kV 500V/19kV 500V/19kV 500V/17.3kV 500V/17.3kV Full drift voltage and potential strips ON in big TPCs, 500V on sense wires, Gating Grid turned on. (BL)
Wed, 11.03.2026 21:41:17 600V/13kV 600V/8.16kV 600V/13kV 600V/17.3kV 600V/19kV 600V/19kV 600V/17.3kV 600V/17.3kV +100V on sense wires, now at 600V (BL)
Wed, 11.03.2026 22:48:02 700V/13kV 700V/8.16kV 700V/13kV 700V/17.3kV 700V/19kV 700V/19kV 700V/17.3kV 700V/17.3kV +100V on sense wires, now at 700V (BL)
Wed, 11.03.2026 23:48:39 800V/13kV 800V/8.16kV 790V/13kV 800V/17.3kV 800V(SR)/740V(HR)/19kV 800V(SR)/740V(HR)/19kV 800V/17.3kV 800V/17.3kV +100V except MTPC HR and VTPC2. MTPC and VTPC2 now at 300V below nominal, going to hold there overnight. (BL)
Thu, 12.03.2026 00:07:56 900V/13kV 850V/8.16kV 790V/13kV 850V/17.3kV 800V(SR)/740V(HR)/19kV 800V(SR)/740V(HR)/19kV 850V/17.3kV 850V/17.3kV +100V on VTPC1, +50V on small TPCs. All TPCs (except VTPC1 sectors 1, 4) are at 300V below nominal (BL)
Thu, 12.03.2026 15:41:21 950V(1,4)/900V/13kV 850V/8.16kV 790V/13kV 850V/17.3kV 800V(SR)/740V(HR)/19kV 800V(SR)/740V(HR)/19kV 850V/17.3kV 850V/17.3kV +50V on VTPC1 sectors 1, 4. Now all are at 300V below nominal (BL)
Thu, 12.03.2026 23:07:44 1050V(1,4)/1000V/13kV 950V/8.16kV 890V/13kV 950V/17.3kV 900V(SR)/840V(HR)/19kV 900V(SR)/840V(HR)/19kV 950V/17.3kV 950V/17.3kV +100V. Now all are at 200V below nominal (BL)
Fri, 13.03.2026 16:40:06 1100V(1,4)/1050V/13kV 1000V/8.16kV 940V/13kV 1000V/17.3kV 950V(SR)/890V(HR)/19kV 950V(SR)/890V(HR)/19kV 1000V/17.3kV 1000V/17.3kV Beginning raise to nominal with +50V to all chambers (BL)
Sat, 14.03.2026 18:46:46 1250V(1,4)/1200V/13kV 1150V/8.16kV 1090V/13kV 1150V/17.3kV 1100V(SR)/1040V(HR)/19kV 1100V(SR)/1040V(HR)/19kV 1150V/17.3kV 1150V/17.3kV Nominal voltage was reached around 23:00 last night. (BL)


Fri, 13.03.2026 17:56:52 At 17:23, began raising GRC voltage as well from its parking value of 1100V, toward nominal 1600V.
Sat, 14.03.2026 18:47:51 During the final raising to nominal, it was discovered that FTPC1 was not actually connected. Had to turn off all FTPC1 voltages plus FTPC2+3 sense wires in order to connect them. FTPC2+3 were raised back to normal relatively quickly, but since FTPC1 had not actually been stressed yet we had to raise it slowly.

Wed, 18.03.2026 07:44:17 In preparation for power cut in EHN1, all TPC voltages and power supplies were turned off.
Wed, 18.03.2026 14:04:54 TPC drift voltage and gating grid were turned on before lunch. Now raising TPCs (and GRC) back to nominal sense wire voltages gradually in steps.
ELOG V3.1.3-7933898